# Scalable Private Search with Wally

### v5

@@ 8 Related Work
Another related area of work is differentially-private multi-server PIR schemes [33, 2]. In both works, servers learn differentially private, noised histograms over queries. The former, [33], has clients generate fake queries whereas the latter, [2], has the servers inject fake queries. Both require a secure multiparty computation over the entire database. Notably, [2] can serve one million queries in about two seconds on a 100K entry database with eight servers in a dishonest majority security model and a ten minute pre-computation. Their offline and online computation grows linearly with the size of the server’s database.
Differential privacy has also been applied to multiple anonymous messaging systems [36, 34, 30]. These are large (many server) mix-net-like protocols where the servers add in fake queries to hide user messages with a differential privacy guarantee. Though they are for a different application, their use of differential privacy is similar to Wally’s.
9 Conclusion
@@ Appendix A Optimizations
A.1 Private Nearest Neighbor Search Optimizations
@@ A.2 PIR Optimizations
In general, the cost of a ciphertext-ciphertext multiplication is much higher than a ciphertext rotation and both are much higher than addition or plaintext multiplication. For example, Table 1 shows ciphertext addition and plaintext multiplication as less than $100\mu s$ while ciphertext rotation is $0.5ms$ and ciphertext multiplication is $2.5ms$.
@@
We apply the lazy rescaling technique of Kim et al [28] in MulPIR’s last step, the inner product between ciphertexts $\mathbf{r}^{T}\mathbf{a}$. The main idea here is that BFV multiplications first multiply polynomials over the integers, then scale and round back to integers modulo $Q$. Kim et al [28, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one. Lazy rescaling yielded a $15-20\%$ improvement in our MulPIR implementation.

### v6

@@ IX Related Work
Differentially-private multi-server (information-theoretic MPC) PIR schemes [26, 27] are not immediately applicable to private search since PIR itself only supports exact matches, but these works’ use of DP is related to our use of DP since they use DP to lower their online computation while leaking information to the servers. In both works, servers learn differentially private, noised histograms of queries over database entries. Note, this compares to Wally where the server learns noised histograms of queries over clusters and only requires one server. The latter [27] is more efficient and requires a multi-server PIR over the entire database once on the database for a batch of queries. Notably, [27] can serve one million queries in about two seconds on a 100K entry database with eight servers in a dishonest majority security model and a ten minute pre-computation. Their offline and online computation grows linearly with the size of the server’s database. Wally, on the other hand, has no offline computation and is sublinear in the database size.
Anonymization networks.
Anonymization networks and mix-nets [64, 24] anonymize the user’s identity but send the query in the clear. Only applying these methods to search systems would not hide users’ private information and are broken by a semi-honest server [51].
Trusted hardware.
Trusted hardware is another tool used in private search systems [65, 66, 67]. However, trusted hardware search systems are only as strong as their underlying hardware [68, 69, 70, 71, 72].
Concurrent works.
Pacmann [43] efficiently reduces graph-based private nearest neighbors search to many PIR queries on a graph. Pacmann uses offline-online PIR which requires the client to stream the database beforehand, then each PIR computation is sublinear since it uses a recent sublinear PIR protocol [73]. As a result, it is a cryptographic, fully-oblivious private search system and offers strong accuracy results from their use of graph-based nearest neighbors search. However, Pacmann requires each client stream the entire database offline whereas Wally requires no offline computation. Further, Pacmann requires many rounds per query, unlike Wally which only requires one round for small database entries and two rounds for large database entries.
X Conclusion
@@ Appendix A Optimizations
A-A Private Nearest Neighbor Search Optimizations
@@ A-B PIR Optimizations
In general, the cost of a ciphertext-ciphertext multiplication is much higher than a ciphertext rotation and both are much higher than addition or plaintext multiplication. For example, Table I shows ciphertext addition and plaintext multiplication as less than $100\mu s$ while ciphertext rotation is $0.5ms$ and ciphertext multiplication is $2.5ms$.
@@
We apply the lazy rescaling technique of Kim et al [39] in MulPIR’s last step, the inner product between ciphertexts $\mathbf{r}^{T}\mathbf{a}$. The main idea here is that BFV multiplications first multiply polynomials over the integers, then scale and round back to integers modulo $Q$. Kim et al [39, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one. Lazy rescaling yielded a $15-20\%$ improvement in our MulPIR implementation.